# KTWIN: A Serverless Kubernetes-based Digital Twin Platform

### v1

@@ Abstract
journal: Internet of Things
@@ 1 Introduction
The growing number of connected devices and the increasing volume of data generated make it challenging to move and process all the data in the Cloud, especially for time-sensitive application requirements. The advances of Mobile Edge Computing (MEC) technologies drove the deployment of micro data centers closer to end users to fit time requirements and, at the same time, overcome computing resource constraints and energy consumption restrictions of the Internet of Things IoT edge devices [2], [3]. In addition, Mobile Cloud Computing (MCC) architecture emerges as a solution for offloading data processing from edge devices to centralized cloud data centers, thus enhancing the user experience and providing greater scalability with access to extensive cloud resources [4]. In such a complex and diverse landscape, building, deploying, and operating Digital Twin components becomes difficult; hence, it is important to establish operations patterns and rely on automation and abstractions to reduce operational burden and costs.
@@
Hence, in this research, we propose KTWIN, a Kubernetes-based platform, that allows DT operators to easily define models through open standards and configure details of the underlying service and infrastructure, such as resource allocation, container-based services deployments, and auto-scaling policies. KTWIN was designed using the state-of-the-art Cloud Native open-source technologies, and it can be deployed in any Kubernetes cluster, in the cloud, or at edge locations. KTWIN leverages the rich environment of Kubernetes and allows for further enhancements and open collaboration within the DT industry and academy. Additionally, KTWIN maps and orchestrates domain entity specification to containers with implementations that can be defined by the DT owners. The benefits of using KTWIN include: (1) reduction of operational costs by automating manual operations, (2) more control and flexibility over the underlying container-based services that compose the DT landscape, (3) can be deployed at any Kubernetes cluster (on-premise, in the edge, in the cloud) reducing the chance of vendor lock-in.
@@
2 Previous Work in Digital Twins
In recent years, the Digital Twin research field has seen significant advancements and contributions from academia leveraging the benefits of DT systems to several industries. Researchers say that the Digital Twin market size can be worth US\$ 110.1 billion by 2028 [6]. This section reviews previous work on Digital Twin systems including applying Digital Twins for different use cases, Digital Twin Architectures, and existing cloud offerings for DT solutions.
@@ 2.1 Digital Twin Use Cases
Digital Twins can improve the energy industry by providing valuable insights into system performance and optimization, reducing power downtime, providing efficient energy planning to fit the supply and demand, as well as optimizing energy household consumption. In this context, Fathy and Jaber [11] propose a data-driven multi-layer Digital Twin of the energy system that aims to mirror the actual energy consumption of households in the form of a Household Digital Twin (HDT). The model intends to improve the efficiency of energy production, modeled as an Energy Production Digital Twin (EDT), by flattening the daily energy demand levels. This is done by collaboratively reorganizing the energy consumption patterns of residential homes to avoid peak demands while accommodating the resident’s needs and reducing their energy costs. In addition, Gao et al. [12] present a Digital Twin-based approach to optimize the operation of an automatic stacking crane handling containers regarding energy consumption. The article developed a virtual container yard that syncs real container yard information to improve automatic stacking crane scheduling. Then, a mathematical model followed by a Q-learning-based optimization step is applied to minimize the total energy consumption for completing all tasks. The output can be used by managers and operators to choose the appropriate strategy to accomplish energy-saving and efficiency goals.
In the Healthcare ecosystem, Digital Twins can be applied for many use cases such as creating a model replica of a single patient for health monitoring and treatment planning, simulating and testing a medical device during the design phase or even modeling an entire healthcare system to monitor to analyze vaccination campaigns, for instance. Pilati et al. [13] developed a Digital Twin which integrates the physical and virtual health systems to create a sustainable and dynamic vaccination center focusing on using the minimum space and resources while guaranteeing a good service for patients. The research implements a discrete event simulation model to try to find the best configuration in terms of human resources and queues for a real vaccination clinic to maximize the number of patients vaccinated in the shortest period.
@@
2.2 Digital Twin Architecture
Talasila et al. [16] propose a Digital Twin framework to manage Digital Twin assets, making them available as a service to other users. The proposed framework automates the management of reusable assets, storage, provision of computing infrastructure, communication and monitoring tasks. Users operate at the level of Digital Twins and delegate the rest of the work to the Digital Twin as a service framework. The paper does not present any real implementation or evaluation scenarios, relying on high-level design aspects of what a Digital Twin platform should contain.
Picone et al. article [18] presents the Edge Digital Twins (EDT) architectural model and its implementation. EDT enables lightweight replication of physical devices and provides a digital abstraction layer to facilitate the interaction with IoT devices and services, enhancing IoT digitalization and interoperability. The EDT architectural blueprint targets to be complementary rather than in competition with existing edge modules and DT solutions (also in the cloud). The proposal provides the following advantages: remove the responsibility from physical assets to handle the integration with cloud DTs, reduce vendor lock-in with cloud providers by introducing EDTs as middleware layers between devices and cloud DT services, and allow applications to interact with EDT locally with MQTT protocol improving performance and reliability.
The Digital Twin literature still has many unexplored study areas, missing an actual deployment of a Digital Twin implementation built on top of Cloud-Native solutions. To the best of our knowledge, no previous work proposes the usage of an event-driven serverless-based architecture in the context of Digital Twin systems, offering serverless abstractions for end-users to combine model design flexibility and system-specific settings. This work aims to fill this gap, providing a comprehensive and scalable solution for Digital Twin research and real-world use case applications.
@@ 2.3 Digital Twins: Existing Services and Frameworks
The existing PaaS solutions offer a wide variety of features, but they do not provide enough flexibility to deploy Digital Twin solutions to different landscapes, such as edge locations, and limit organizations to work with hybrid or multi-cloud environments. Moreover, they mostly are vendor-specific implementations, so end-users cannot move their solutions to another vendor stack without a considerable amount of rework. Lastly, all the services mentioned above offer less control over operations and infrastructure, preventing users from personalizing settings, and monitoring application health and events propagation. In this context, an open-source Digital Twin platform built with Cloud Native solutions can offer flexibility and independence to organizations to build microservice-based distributed systems to leverage more Digital Twin use cases.
@@ 3 KTWIN Architecture Design
KTWIN, a Kubernetes-based Serverless platform for Digital Twins, automates the orchestration of Digital Twin services allowing operators to easily define twin’s data models and their relationships reducing the complexity of deploying and managing DT services. The KTWIN architecture design requirements methodology definition considered the Digital Twin Capability Periodic Table framework defined by the Digital Twin Consortium [23], a collaborative partnership with industry, academia and government experts, that aims to drive the awareness, adoption, interoperability, and development of Digital Twin technology. The actual design and implementation fulfills some of the requirements established by the group focusing on the system-design aspects required for a Digital Twin Serverless platform. Among the fulfilled requirements are Data Acquisition and Injection, Real-time Processing, IoT Systems Integration, Temporal Data Store, DT Instance Repository, System Monitoring, Reliability and Resilience. The following subsections present KTWIN main components, the implemented custom resources and eventing routing mechanisms, and aspects regarding data management and storage.
@@ 4.1 KTWIN Architecture Implementation
The main personas involved in KTWIN development and operational process are: DT Owner, DT Developer and DT Operator. The DT Operator creates the Custom Resource Definition (CRD) files to model the Digital Twin use case within KTWIN and apply the resources using Kubernetes CLI. The CRD files are based on the ontology definition originally provided by the DT Owner. The provided files are processed by KTWIN Operator, developed in Golang using Kubebuilder, an open-source SDK library that facilitates seamless interaction with Kubernetes Control Plane APIs. These CRD files are stored in the Kubernetes etdc storage, ensuring reliable and consistent data management. The KTWIN Operator creates the required Application Plane resources to instantiate the defined use case. The container-based functions implemented by the DT Developer are pulled from a container registry and deployed as a Twin Service. Finally, the DT Owner can analyze the processed data using some visualization tool that fetches the data from the Event Store.
@@ 4.2 Kubernetes Custom Resources
A Kubernetes Custom Resource Definition (CRD) is a way to extend the Kubernetes API by defining and managing custom resource types. KTWIN makes use of CRD to allow Digital Twin owners to define Digital Twin entities, their attributes and relation, in addition to system-specific definitions, such as allocated CPU or memory and auto-scaling policies. The developed KTWIN prototype enhances the open-source Digital Twin Definition language (DTDL) by adding system-specific configuration to the ontology-based definition language. The implemented prototype has two DTDL-based CRDs: Twin Interface and Twin Instance.

### v3

@@ Abstract
journal: Computer Networks
@@ 1 Introduction
The growing number of connected devices and the increasing volume of data generated make it challenging to move and process all the data in the Cloud, especially for time-sensitive application requirements. The advances of Mobile Edge Computing (MEC) technologies drove the deployment of micro data centers closer to end users to fit time requirements and, at the same time, overcome computing resource constraints and energy consumption restrictions of the Internet of Things IoT edge devices [2]. In addition, Mobile Cloud Computing (MCC) architecture emerges as a solution for offloading data processing from edge devices to centralized cloud data centers, thus enhancing the user experience and providing greater scalability with access to extensive cloud resources [3]. In such a complex and diverse landscape, building, deploying, and operating Digital Twin components becomes difficult; hence, it is important to establish operations patterns and rely on automation and abstractions to reduce operational burden and costs.
@@
Hence, in this research, we propose KTWIN, a Serverless Kubernetes-based platform, that allows DT operators to easily define models through open standards and configure details of the underlying service and infrastructure, such as resource allocation, container-based services deployments, and auto-scaling policies. KTWIN utilizes Serverless principles to abstract infrastructure management, enabling on-demand scaling, efficient resource utilization, and reduced operational overhead while ensuring high availability and responsiveness for Digital Twin applications. The platform was designed using the state-of-the-art Cloud Native open-source technologies, and it can be deployed in any Kubernetes cluster, in the cloud, or at edge locations. KTWIN leverages the rich environment of Kubernetes and allows for further enhancements and open collaboration within the DT industry and academy. Additionally, KTWIN maps and orchestrates domain entity specification to containers with implementations that can be defined by the DT owners. The benefits of using KTWIN include: (1) reduction of operational costs by automating manual operations, (2) more control and flexibility over the underlying container-based services that compose the DT landscape, (3) can be deployed at any Kubernetes cluster (on-premise, in the edge, in the cloud) reducing the chance of vendor lock-in.
@@
2 Background
In recent years, the Digital Twin (DT) research field has seen significant advancements and contributions from academia leveraging the benefits of DT systems to several industries. Researchers say that the Digital Twin market size can be worth US\$ 110.1 billion by 2028 [5]. This section reviews the different use cases that Digital Twins can be applied, as well as the existing platforms and frameworks offerings and their limitations.
@@ 2.1 Digital Twin Use Cases
A Digital Twin is a virtual representation of a physical system or process, designed to facilitate bidirectional communication between the real-world system and its digital counterpart [6]. DTs enable the development of value-added services, such as real-time monitoring, predictive maintenance, optimization, and simulation, that can be adopted by a variety of industries such as Manufacturing, Smart Cities, Energy and Healthcare.
@@
Digital Twins can improve the energy industry by providing valuable insights into system performance and optimization, reducing power downtime, providing efficient energy planning to fit the supply and demand, as well as optimizing energy household consumption. In this context, Fathy and Jaber [11] propose a data-driven multi-layer Digital Twin of the energy system that aims to mirror the actual energy consumption of households in the form of a Household Digital Twin (HDT). The model intends to improve the efficiency of energy production, modeled as an Energy Production Digital Twin (EDT), by flattening the daily energy demand levels. This is done by collaboratively reorganizing the energy consumption patterns of residential homes to avoid peak demands while accommodating the resident’s needs and reducing their energy costs. In addition, Gao et al. [12] present a Digital Twin-based approach to optimize the operation of an automatic stacking crane handling containers regarding energy consumption. The article developed a virtual container yard that syncs real container yard information to improve automatic stacking crane scheduling. Then, a mathematical model followed by a Q-learning-based optimization step is applied to minimize the total energy consumption for completing all tasks. The output can be used by managers and operators to choose the appropriate strategy to accomplish energy-saving and efficiency goals. In the Healthcare ecosystem, Digital Twins can be applied for many use cases such as creating a model replica of a single patient for health monitoring and treatment planning, simulating and testing a medical device during the design phase or even modeling an entire healthcare system to monitor to analyze vaccination campaigns, for instance. Pilati et al. [13] developed a Digital Twin which integrates the physical and virtual health systems to create a sustainable and dynamic vaccination center focusing on using the minimum space and resources while guaranteeing a good service for patients. The research implements a discrete event simulation model to try to find the best configuration in terms of human resources and queues for a real vaccination clinic to maximize the number of patients vaccinated in the shortest period.
@@ 2.2 Digital Twins: Existing Services and Frameworks
The existing PaaS solutions offer a wide variety of features, but they do not provide enough flexibility to deploy Digital Twin solutions to different landscapes, such as edge locations, and limit organizations to work with hybrid or multi-cloud environments. Moreover, they mostly are vendor-specific implementations, reducing interoperability and reusability [6], and limiting organizations to move their solutions to another vendor stack without a considerable amount of rework. Lastly, all the services mentioned above offer less control over operations and infrastructure, preventing users from personalizing settings, and monitoring application health and events propagation. In this context, an open-source Digital Twin platform built with Cloud Native solutions can offer flexibility and independence to organizations to build microservice-based distributed systems to leverage more Digital Twin use cases.
3 Previous Work in Digital Twins
This section delves into the state of the art in DT research, highlighting recent advancements, frameworks, and architectures, in the field. Grieves et al.[18] formalized that a DT system is a composition of three primary elements: a real-world space composed of a set of objects or systems, a virtual space representation of the real-world space, and a link in which the data flow between both spaces. Tao et al. [19] enhanced the definition and established the Five-Dimensional Digital Twin model composed by: a physical entity model, virtual entity model, services model, data model and connection model.
Several prior works have proposed frameworks for implementing DTs). Talasila et al. [20] propose a Digital Twin framework to manage Digital Twin assets, making them available as a service to other users. The proposed framework automates the management of reusable assets, storage, provision of computing infrastructure, communication and monitoring tasks. Users operate at the level of Digital Twins and delegate the rest of the work to the Digital Twin as a service framework. The paper does not present any real implementation or evaluation scenarios, relying on high-level design aspects of what a Digital Twin platform should contain.
Picone et al. article [21] presents the Edge Digital Twins (EDT) architectural model and its implementation. EDT enables lightweight replication of physical devices and provides a digital abstraction layer to facilitate the interaction with IoT devices and services, enhancing IoT digitalization and interoperability. The EDT architectural blackprint targets to be complementary rather than in competition with existing edge modules and DT solutions (also in the cloud). The proposal provides the following advantages: remove the responsibility from physical assets to handle the integration with cloud DTs, reduce vendor lock-in with cloud providers by introducing EDTs as middleware layers between devices and cloud DT services, and allow applications to interact with EDT locally with MQTT protocol improving performance and reliability.
Context-aware and adaptive Digital Twin models have been studied by researches. Hribernik et al. [23] discusses the potential of context-aware, autonomous, and adaptive DTs being the building blocks for tomorrow’s Digital Factories. Bellavista et al. [24] propose a vision of DT design and implementation to contain adaptive, autonomous, and context-aware functionalities. The paper also propose and implements a set of reusable micro-services design patterns allowing engineers to meet these new demanding requirements while keeping complexity and management costs under control.
Minerva et al. [25] provide a comprehensive survey of technologies, scenarios, application cases, and architectural models relevant to the implementation of the DT concept. The study consolidates key definitions, specifications, and implementations of DTs across diverse technological domains, offering insight into how the concept has evolved and been adopted in various industries. The article explores the applicability of DTs in key IoT application scenarios and proposes a framework for evaluating its implementation with modern software architectures.
Bicocchi et al. [26] explored the concept of end-to-end trustworthiness in the context of DTs. The research identifies the six key characteristics that enable trustworthiness, i.e., entanglement awareness, variable load resilience, edge-to-cloud continuum mobility, declarative representation, observability, and security. In addition, The authors designed a blackprint architecture that natively supports end-to-end trustworthiness and discussed a proof-of-concept implementation.
The Digital Twin literature still has many unexplored study areas, missing an actual deployment of a Digital Twin implementation built on top of Cloud-Native solutions. To the best of our knowledge, no prior work has proposed an event-driven, Serverless-based architecture for DT systems that provides Serverless abstractions to enable end-users to seamlessly combine flexible model design with system-specific configurations. This work addresses these gaps by introducing a novel solution that includes (1) an extended DTDL for higher-level abstraction, enabling end users to define domain entities while maintaining control over underlying services, (2) an automated process to orchestrate DT components creation, modification and deletion, including messaging, routing, and data storage, and (3) the development and deploiyment of a Smart City DT use case to validate the platform. This novel approach is expected to significantly enhance the adoption of DT technology by simplifying the development and deployment of new and diverse use cases.
@@ 4 KTWIN Architecture Design
KTWIN, a Kubernetes-based Serverless platform for Digital Twins, automates the orchestration of Digital Twin services allowing operators to easily define twin’s data models and their relationships reducing the complexity of deploying and managing DT systems. The following subsections outline the architectural requirements and decisions of KTWIN and detail its main components. Additionally, they describe the implemented custom resources, event routing mechanisms, and key aspects of data management and storage.
4.1 Architectural Requirements and Design Choices
Understanding the distinct needs of the personas involved in the design, development, and operational lifecycle of Digital Twins is instrumental in shaping KTWIN’s architectural requirements and in ensuring that the platform is tailored to meet their specific challenges and objectives. Therefore, we defined KTWIN revolving around three primary personas, described as follows:
  • 1.
    The DT Domain Expert is primarily responsible for the strategic direction and governance of the Digital Twin ecosystem. Their responsibilities include defining the functional requirements by modeling entities and describing their behaviors, often leveraging an ontology-based language to ensure precise representation of domain-specific entities. This requires a system capable of adapting to diverse domains through open and flexible modeling tools.
  • 2.
    The DT Developer focuses on implementing and validating services that encapsulate the domain logic defined by the DT Domain Expert. The developer seeks for abstractions such as Function as a Service (FaaS), service containerization, and the usage of Software Development Kits (SDKs), to streamline service creation and deployment.
  • 3.
    Lastly, the DT Operator responsibilities include the deployment and maintenance of the Digital Twin solution, combining domain requirements with system-level configurations. Their role requires an unified, vendor-neutral specification language that enables seamless deployment across cloud, on-premise, and edge environments.
4.1.1 Architectural Requirements
Orchestration Abstraction. The orchestration abstraction provided by KTWIN significantly streamlines the process of designing and managing Digital Twin systems by reducing complexity in implementation and maintenance. High-level data modeling, powered by ontology-based and open-standard tools, allows users to define and manage entities, relationships, and behaviors in a structured and reusable manner. This approach ensures cross-domain compatibility, enabling a consistent representation of Digital Twin systems. Furthermore, the simplified configuration process enables users to have an enhanced control and flexibility over the underlying services, combining domain-specific configurations with system-specific requirements to set up and operate the platform effectively. By abstracting underlying complexities, KTWIN democratizes access to sophisticated Digital Twin capabilities, reducing barriers to adoption and minimizing implementation and maintenance costs.
Self-Contained Modules. The modularity inherent in the platform ensures that the developers define twin behaviors as self-contained, interchangeable functions, allowing development, deployment, updates, and scaling of the individual services. By defining twin’s implementation as independent and self-contained functions, KTWIN enhances maintainability, as updates or replacements to a module can be performed with minimal disruption. Additionally, modularity fosters reusability, allowing components to be repurposed across multiple Digital Twin implementations, reducing redundancy and accelerating development cycles.
Data Acquisition and Ingestion. Data acquisition and ingestion in KTWIN are critical for integrating real-world data into Digital Twin systems, ensuring accurate and up-to-date models. The platform supports diverse data sources, including IoT devices and edge systems, through a centralized broker that supports multiple protocols such as MQTT, HTTP, and AMQP. This capability allows KTWIN to collect, route, and process events efficiently, whether they originate from high-frequency sensors or intermittent data sources. Furthermore, the system can adapt to increasing data volumes and evolving workload demands effectively, allowing individual components to scale independently based on real-time demand. The event-driven design also ensures resilience, as services can continue functioning even during high loads or partial failures.
Services Interoperability. Interoperability enables seamless integration and operation across diverse systems and environments. By adhering to open standards and leveraging cloud-native tools, the platform ensures compatibility with a broad range of technologies, promoting portability and reducing vendor lock-in. Additionally, KTWIN supports deployment across multi-cloud, edge, and on-premises infrastructures, allowing organizations to tailor the platform to their specific operational needs. This capability ensures that Digital Twin systems can operate efficiently regardless of the underlying environment, making KTWIN a versatile solution for heterogeneous ecosystems.
The aforementioned requirements provide a clear understanding of the key architectural considerations that shaped the design and development of KTWIN platform. They played a pivotal role in guiding the architectural decision-making process, detailed as follows.
4.1.2 Architectural Design Choices
Serverless. KTWIN-based applications requires high scalability, low operational overhead, and the ability to handle variable workloads efficiently. Serverless computing platforms provide Function(s)-as-a-Service (FaaS) by hosting individual callable functions. These platforms provide reduced hosting costs, fast deployment, automatic scaling, high availability, fault tolerance, and dynamic elasticity through automatic provisioning and management of compute infrastructure [29], making it ideal for handling variable workloads with low operational overhead. The proposed KTWIN Serverless architecture implements automated services infrastructure provisioning based on the provided ontology resources definition. The usage of vendor-specific Serverless platforms come with certain limitations, including cold start latency, execution time restrictions, and the risk of vendor lock-in [29].
Event-Driven. Event-Driven Architecture (EDA) allows to decouple services, scale effectively, and process events in real-time. KTWIN employs a centralized broker ensuring the processing of events regardless of their source and supporting multiple communications protocols, including MQTT, HTTP and AMQP. Additionally, its event-based routing mechanism is automatically configured based on ontology’s relationship definitions, simplifying development and operation efforts for the platform users. The KTWIN event-driven architecture provides the flexibility and resilience necessary for Digital Twin systems, allowing independent service evolution, fault tolerance, and improved extensibility. Furthermore, the event-driven microservices design seamlessly integrates with Serverless architectures, offering inherent benefits such as high availability, fault resilience, and cost efficiency.
Kubernetes and Container-based Services. Kubernetes serves as a backbone for KTWIN, providing scalability, flexibility, and resilience to manage complex containerized applications across diverse environments. KTWIN leverages on Kubernetes operators and Custom Resource Definitions (CRDs) to automate the management and provisioning of cluster infrastructure. Kubernetes also enables multi-cloud and edge-based deployments strategies needed for microservice-based architecture while avoiding vendor lock-in. Lastly, Kubernetes integrates with rich cloud-native environments, including Serverless platforms. This integration enables KTWIN to leverage Function-as-a-Service (FaaS) models.
Cloud Native-based Open Platform. KTWIN leverages the rich Cloud Native ecosystem to meet scalability, flexibility, and resilience needs. The adoption of open solutions and standards benefits from an open, innovative and vendor-neutral ecosystem that encourages innovation. The rich open ecosystem also provides cost-efficient and mature tools to manage deployment, observability and security. In terms of ontology-based domain definition languages, KTWIN leverages the Digital Twins Definition Language (DTDL) to offer a flexible and open approach for describing digital twin models of smart devices.
KTWIN advances the state-of-the-art on DT platforms by addressing key limitations in existing approaches while introducing innovative capabilities that have yet to be thoroughly explored in the literature. It achieves this by (1) reducing operational costs through the automation of manual processes, streamlining deployment and management tasks, (2) offering enhanced control and flexibility over the container-based services that constitute the DT landscape, enabling tailored configurations to meet diverse requirements for DTs of any domain, and (3) supporting deployment across any Kubernetes cluster, whether on-premises, at the edge, or in the cloud, minimizing the risk of vendor lock-in, and providing organizations with greater autonomy and adaptability.
@@ 5.1 KTWIN Architecture Implementation
The main personas involved in KTWIN development and operational process are: DT Domain Expert, DT Developer and DT Operator. The DT Operator creates the Custom Resource Definition (CRD) files to model the Digital Twin use case within KTWIN and apply the resources using Kubernetes CLI. The CRD files are based on the ontology definition originally provided by the DT Domain Expert. The provided files are processed by KTWIN Operator, developed in Golang using Kubebuilder, an open-source SDK library that facilitates seamless interaction with Kubernetes Control Plane APIs. These CRD files are stored in Kubernetes etdc storage, ensuring reliable and consistent data management. The KTWIN Operator creates the required Application Plane resources to instantiate the defined use case. The container-based functions implemented by the DT Developer are pulled from a container registry and deployed as a Twin Service. Finally, the DT Domain Expert can analyze the processed data using some visualization tool that fetches the data from the Event Store.
@@ 5.2 Kubernetes Custom Resources
A Kubernetes Custom Resource Definition (CRD) is a way to extend the Kubernetes API by defining and managing custom resource types. KTWIN makes use of CRD to allow Digital Twin Domain Experts to define Digital Twin entities, their attributes and relation, in addition to system-specific definitions, such as allocated CPU or memory and auto-scaling policies. The developed KTWIN prototype enhances the open-source Digital Twin Definition language (DTDL) by adding system-specific configuration to the ontology-based definition language. The implemented prototype has two DTDL-based CRDs: Twin Interface and Twin Instance.